# Chasing the Timber Trail: Machine Learning to Reveal Harvest Location Misrepresentation

### v1

@@ Abstract.
Illegal logging poses a significant threat to global biodiversity, climate stability, and depresses international prices for legal wood harvesting and responsible forest products trade, affecting livelihoods and communities across the globe. Stable isotope ratio analysis (SIRA) is rapidly becoming an important tool for determining the harvest location of traded, organic, products. The spatial pattern in stable isotope ratio values depends on factors such as atmospheric and environmental conditions and can thus be used for geographical identification. We present here the results of a deployed machine learning pipeline where we leverage both isotope values and atmospheric variables to determine timber harvest location. Additionally, the pipeline incorporates uncertainty estimation to facilitate the interpretation of harvest location determination for analysts. We present our experiments on a collection of oak (Quercus spp.) tree samples from its global range. Our pipeline outperforms comparable state-of-the-art models determining geographic harvest origin of commercially traded wood products, and has been used by European enforcement agencies to identify illicit Russian and Belarusian timber entering the EU market. We also identify opportunities for further advancement of our framework and how it can be generalized to help identify the origin of falsely labeled organic products throughout the supply chain.
@@ 1. Introduction
Illegal logging is a global problem with serious economic, social, and ecological impacts. It is the most profitable natural resource crime with an estimated annual value between \$52 billion and \$157 billion, as well as the third largest transnational crime, after counterfeiting and drug trafficking (tra, [n. d.]). Illegal logging causes destruction to forest ecosystems, the local flora and fauna, and the people whose livelihoods depend upon forests for goods and services. The effects of illegal logging are not only felt in the area where the illicit activity takes place but also impact associated trade and international markets. Two decades ago it was determined that the value of US-harvested hardwoods (e.g., maple, oak, ash, poplar) had been depressed by 7-16% due to the amount of illegal timber in global trade (Seneca Creek Associates, 2004). Given the complexity of global supply chains and the environmental and societal harm that practices such as illegal resource harvesting present, it is increasingly difficult for stakeholders (e.g., consumers, businesses involved in international trade, and governments seeking to enforce regulations) to ensure that products do not contribute to such environmental and social ills.
@@
  • In 2015, US flooring importer Lumber Liquidators plead guilty to environmental crimes related to importing illegally harvested Russian wood, which included felony misdeclarations on its US import documentation, claiming the species of oak used was harvested in Germany, even though that specific species only grows in Northeast Asia (off, 2015);
  • Due to Russia’s 2022 war in Ukraine and the sanctions regime put in place by many countries, there has been increased scrutiny on sanctions evasion of Russian timber entering Europe by misdeclaring the harvest origin (Earthsight, [n. d.]).
What is now possible, due to global efforts including the effort described here, is to use stable isotope ratio analysis (SIRA) to validate the claimed geographic harvest origins from a finished forest product such as furniture or flooring to the forests where the wood was harvested. Stable isotopes are chemical variants of elements that do not go through radioactive decay. By measuring stable isotope values through mass spectrometry (Barrie and Prosser, 1996) and comparing it to a location-referenced database, we can help understand the origin of organic products. Previous work has shown how it can be used to trace the origin of items such as timber, seafood, agricultural products, and fiber such as cotton (Truszkowski et al., 2025; Mortier et al., 2024; Watkinson et al., 2022; Cusa et al., 2022; Wang et al., 2020; Meier-Augenstein et al., 2014).
We focus on wood and forest products as they play a critical role in the fight against climate change and, as mentioned earlier, the illegal harvesting of trees is the most profitable natural resource crime (tra, [n. d.]; Messetchkova, 2021). The use of stable isotope testing as a tool to determine the origin of such products is seen as a critical mechanism to identify instances of false or misrepresented locations of harvest claims and of wood and forest products, thus assisting to uncover instances of illegal logging and sanctions evasion.
@@
    Our work is a deployed machine learning pipeline wherein physical samples are collected, subject to tests, and injected into our model to help European enforcement agencies in combating illegal timber trade. Notably, our framework has been rigorously evaluated post-deployment and demonstrated to be instrumental in identifying illicit Russian and Belarusian timber entering the EU market; see the coverage of our work and enforcement actions from the New York Times and Brussels Times at (Nazaryan, 2024; Times, [n. d.]).
@@ 2.2. Atmospheric Data Layers
Stable isotope ratios are largely dependent on various environmental factors, including but not limited to precipitation, water vapor pressure, reflected short wave radiation(Vystavna et al., 2021; Elbeltagi et al., 2023), and the ecological process that facilitates the enrichment, or lack thereof, for stable isotopes as observed in oak trees is a lengthy process that can be captured by climatic patterns over a long time. We collect comprehensive data about 25 different atmospheric properties (NASA, 2024; Huffman et al., 2020; Bowen and Revenaugh, 2003) for over 20 years across the world to capture the overall climatic conditions including regularities and anomalies.
@@ 3.3. RQ1: Baseline Experiments
  • Watkinson et al. (2022): This work uses a GPR to predict the isotope ratio values in Quercus spp.
@@
The experiments as detailed in Table 1 demonstrate that, in terms of $R^{2}$, our proposed frameworks GB-SIRA and MTG-SIRA outperform other approaches for 2 out of the 4 isotope ratio prediction tasks. However, even for $\delta^{13}\mathrm{C}$ and $\delta^{34}\mathrm{S}$, the difference w.r.t. the best-performing model is negligible. For isotopes such as $\delta^{2}\mathrm{H}$ and $\delta^{18}\mathrm{O}$, there is significant improvement witnessed.
In terms of RMSE, we also observe the lowest RMSE value for MTG-SIRA for all stable isotope ratios except $\delta^{34}\mathrm{S}$ for which GB-SIRA performs the best. (It is important to note that for Watkinson et al. (2020), RMSE was not reported.) The performance of both GB-SIRA and MTG-SIRA overall compares favorably, demonstrating the effectiveness of the proposed ML framework. In general, the performance for $\delta^{34}\mathrm{S}$ and $\delta^{13}\mathrm{C}$ is not high across the models, a result in line with past related work. Finally, the overall performance improvement on the multitask model MTG-SIRA indicates the effectiveness of modeling inter-task dependency.
@@ 3.7. RQ5: Evaluating Claims of Timber Origin
As mentioned earlier, our approach is in use by European enforcement agencies and has been shown to be effective in evaluating claims. However, because we cannot showcase results on real data, we formulate a controlled experiment wherein we modify the true (ground truth) location of a sample by a specified distance $d$. Thus $x$ is modified, resulting in a modified atmospheric vector for $x$. This can be formulated as:
@@ 4. Description of Deployment
Our framework has been rigorously evaluated post-deployment and demonstrated to be instrumental in identifying illicit Russian and Belarusian timber entering the EU market; see the coverage of our work and enforcement actions from the New York Times and Brussels Times at (Nazaryan, 2024; Times, [n. d.]). Due to confidentiality reasons, we present results here on a different species (Oak; Quercus spp.). For instance, see isoscape predictions around the Russia/Ukraine border region in Fig. 5.
@@ 5. Conclusion
We have presented a novel ML framework that uses SIRA to determine the origin of organic products. It serves as a holistic end-to-end predictive pipeline that can facilitate feature selection and systematic analysis of recorded (and predicted) isotope ratio values. Our approach to multimodal feature integration also enables us to leverage the vast information present in atmospheric data by systematic feature selection methods. Further, such incorporation of atmospheric variables into the methodology supports interpretability that prior methods do not support.
@@
Using our methodology to infer the origins of an unseen sample with varying degrees of confidence has proven to be an effective tool used to assist in the enforcement of trade and sustainability regulations governing commodities like timber, such as identification of sanctioned Russian wood and species listed on the Convention on the International Trade in Endangered Species (Mortier et al., 2024; Grove and Rutherford, 2023).
There is promise for additional social and environmental impact of this work, given that the application of our SIRA methods can be applied to origin testing for a variety of organic products. Our work is particularly relevant given the EU’s regulation on deforestation free products and forest-risk commodities (EUDR) which requires proving that products made from forest-risk commodities, e.g., cattle, wood, cocoa, soy, coffee, palm oil, and rubber, do not originate from recently deforested land or have contributed to forest degradation (European Commission, 2023).

### v2

@@ Abstract.
Illegal logging poses a significant threat to global biodiversity, climate stability, and depresses international prices for legal wood harvesting and responsible forest products trade, affecting livelihoods and communities across the globe. Stable isotope ratio analysis (SIRA) is rapidly becoming an important tool for determining the harvest location of traded, organic, products. The spatial pattern in stable isotope ratio values depends on factors such as atmospheric and environmental conditions and can thus be used for geographic origin identification. We present here the results of a deployed machine learning pipeline where we leverage both isotope values and atmospheric variables to determine timber harvest location. Additionally, the pipeline incorporates uncertainty estimation to facilitate the interpretation of harvest location determination for analysts. We present our experiments on a collection of oak (Quercus spp.) tree samples from its global range. Our pipeline outperforms comparable state-of-the-art models determining geographic harvest origin of commercially traded wood products, and has been used by European enforcement agencies to identify harvest location misrepresentation. We also identify opportunities for further advancement of our framework and how it can be generalized to help identify the origin of falsely labeled organic products throughout the supply chain.
@@ 1. Introduction
Illegal logging is a global problem with serious economic, social, and ecological impacts. It is the most profitable natural resource crime with an estimated annual value between \$52 billion and \$157 billion, as well as the third largest transnational crime, after counterfeiting and drug trafficking (Integrity, 2017). Illegal logging causes destruction to forest ecosystems, the local flora and fauna, and the people whose livelihoods depend upon forests for goods and services. The effects of illegal logging are not only felt in the area where the illicit activity takes place but also impact associated trade and international markets. Two decades ago it was determined that the value of US-harvested hardwoods (e.g., maple, oak, ash, poplar) had been depressed by 7-16% due to the amount of illegal timber in global trade (Seneca Creek Associates, LLC., 2004). Given the complexity of global supply chains and the environmental and societal harm that practices such as illegal resource harvesting present, it is increasingly difficult for stakeholders (e.g., consumers, businesses involved in international trade, and governments seeking to enforce regulations) to ensure that products do not contribute to such environmental and social ills.
@@
  • In 2015, US flooring importer Lumber Liquidators plead guilty to environmental crimes related to importing illegally harvested Russian wood, which included felony misdeclarations on its US import documentation, claiming the species of oak used was harvested in Germany, even though that specific species only grows in Northeast Asia (of Justice, 2015);
  • Due to Russia’s 2022 war in Ukraine and the sanctions regime put in place by many countries, there has been increased scrutiny on sanctions evasion of Russian timber entering Europe by misdeclaring the harvest origin (Earthsight, 2025; Times, 2024).
What is now possible, due to global efforts including the effort described here, is to use stable isotope ratio analysis (SIRA) to validate the claimed geographic harvest origins from finished forest products such as furniture or flooring. Stable isotopes are chemical variants of elements that do not go through radioactive decay. By measuring stable isotope values through mass spectrometry (Barrie and Prosser, 1996) and comparing it to a location-referenced database, we can help understand the origin of organic products. Previous work has shown how it can be used to trace the origin of items such as wood, seafood, agricultural products, and cotton (Truszkowski et al., 2025; Mortier et al., 2024; Watkinson et al., 2022; Cusa et al., 2022; Wang et al., 2020; Meier-Augenstein et al., 2014).
We focus on wood and forest products, which play a critical role in the fight against climate change and, as mentioned earlier, the illegal harvesting of trees is the most profitable natural resource crime (Integrity, 2017; Messetchkova, 2021). The use of stable isotope testing as a tool to determine the origin of such products is seen as a critical mechanism to identify instances of false or misrepresented harvest location claims of wood and forest products, thus assisting to uncover instances of illegal logging and sanctions evasion.
@@
    Our work is a deployed machine learning pipeline wherein physical samples are collected, subject to tests, and injected into our model to help European enforcement agencies in combating illegal timber trade by demonstrating that a claimed harvest location other than Russia is not viable. See coverage of our work from the New York Times (Nazaryan, 2024).
@@ 2.2. Atmospheric Data Layers
Stable isotope ratios are largely dependent on various environmental factors, including but not limited to precipitation, water vapor pressure, and reflected shortwave radiation (Vystavna et al., 2021; Elbeltagi et al., 2023). The ecological process that facilitates enrichment, or the lack thereof, for stable isotopes in oak trees is a lengthy process. This process can be influenced by climatic patterns observed over long periods. To capture overall climatic conditions, including regularities and anomalies, we collect comprehensive data on 25 different atmospheric properties (NASA, 2024; Huffman et al., 2020; Bowen and Revenaugh, 2003) spanning over 20 years across the world.
@@ 3.3. RQ1: Baseline Experiments
  • Watkinson et al. (2022): This work uses Universal Kriging to predict the isotope ratio values of timber in Solomon Islands.
@@
The experiments as detailed in Table 1 demonstrate that, in terms of $R^{2}$, our proposed frameworks GB-SIRA and MTG-SIRA outperform other approaches for 2 out of the 4 isotope ratio prediction tasks. However, even for $\delta^{13}\mathrm{C}$ and $\delta^{34}\mathrm{S}$, the difference with respect to the best-performing model is negligible. For isotopes such as $\delta^{2}\mathrm{H}$ and $\delta^{18}\mathrm{O}$, there is significant improvement witnessed.
In terms of RMSE, we also observe the lowest RMSE value for MTG-SIRA for all stable isotope ratios except $\delta^{34}\mathrm{S}$ for which GB-SIRA performs the best. We note that Watkinson et al. (2020) did not report RMSE. The performance of both GB-SIRA and MTG-SIRA overall compares favorably, demonstrating the effectiveness of the proposed ML framework. In general, the performance for $\delta^{34}\mathrm{S}$ and $\delta^{13}\mathrm{C}$ is not high across the models, a result in line with past related work. Finally, the overall performance improvement on the multitask model MTG-SIRA indicates the effectiveness of modeling inter-task dependency.
@@ 3.7. RQ5: Evaluating Claims of Timber Origin
As mentioned earlier, our approach is in use by European enforcement agencies. However, because we cannot showcase results on real data, we formulate a controlled experiment wherein we modify the true (ground truth) location of a sample by a specified distance $d$. Thus $x$ is modified, resulting in a modified atmospheric vector for $x$. This can be formulated as:
@@ 4. Description of Deployment
Our framework is in use by European enforcement agencies to assist in demonstrating that a claimed harvest location other than Russia is not viable. See coverage of our work from the New York Times (Nazaryan, 2024). Due to confidentiality reasons, we present results here on a different species (Oak; Quercus spp.). For instance, see isoscape predictions around the Russia/Ukraine border region in Fig. 5.
@@ 5. Conclusion
We have presented a novel ML framework that uses SIRA to determine the origin of organic products. It serves as a holistic end-to-end predictive pipeline that can facilitate feature selection and systematic analysis of recorded and predicted isotope ratio values. Our approach to multimodal feature integration also enables us to leverage the vast information present in atmospheric data by systematic feature selection methods. Further, such incorporation of atmospheric variables into the methodology supports interpretability that prior methods do not support.
@@
Using our methodology, being able to infer the origins of an unseen sample with varying degrees of confidence has proven to be an effective tool used to enforce timber trade regulations and sanctions, such as the sanction on importing Russian and Belarusian timber in the EU (Mortier et al., 2024; Truszkowski et al., 2025).
There is promise for additional social and environmental impact of this work, given that the application of our SIRA methods can be applied to origin testing for a variety of organic products. Our work is particularly relevant given the EU’s Regulation on Deforestation-free Products (EUDR) which requires proving that products made from forest-risk commodities, e.g., cattle, wood, cocoa, soy, coffee, palm oil, and rubber, do not originate from recently deforested land or have contributed to forest degradation (European Commission, 2023).